# Revealing the Inherent Instructability of Pre-Trained Language Models

### v2

@@ 7 Conclusion
We hypothesize that the pre-training stage enables LLMs to develop the ability to process instructions. To test this, we propose Response Tuning (RT), a method that removes instructions from IT and focuses solely on learning the response distribution. Our extensive experiments demonstrate that RT models—trained only on responses without paired instructions—effectively respond to a wide range of user queries. Moreover, we observe that they can reject unsafe requests by comprehending instructions and invoking learned conditions of refusal. These observations also hold in an in-context learning setting. Together, these results show that establishing an adequate output distribution alone can yield instructable models, supporting our hypothesis that pre-trained models possess the ability to process instructions.
@@
The primary aim of our work is to provide insights into how LLMs shift from pre-trained models to instructable assistants. In our study, RT serves as a tool—intentionally designed as a simplified ablation of IT—but as a standalone method, it exhibits limited performance compared to IT. Future work could build on the insights gained from RT to develop more effective and efficient tuning methods. Moreover, future research could explore the effect of pre-training on more nuanced aspects of instruction-following—such as mitigating sycophancy or social bias (Perez et al., 2023; Sharma et al., 2024).
@@
Our study involves human evaluations to evaluate instruction-following LLMs. The evaluators were hired in compliance with local laws and were paid appropriate compensation. The authors manually reviewed the LLM responses flagged by OpenAI moderation API and confirmed that these pose no harm to human evaluators. In addition, evaluators had the right to immediately stop the evaluation if they wished, and were encouraged to discuss any discomfort with the authors. While we publicly release the codes for safety evaluations, we decide not to release the refusal judge validation set to prevent potential misuse of unsafe or illegal information.
Reproducibility Statement
The training setup of the models including hyperparameters can be found in Section 4.1. We detail the evaluation setup in Appendix A.
@@ Acknowledgements
Seokhyun An is supported by the Korea Presidential Science Scholarship. This work was supported by grants from the Institute of Information & Communications Technology Planning & Evaluation (IITP), funded by the Ministry of Science and ICT (MSIT) of the Korean government: No. RS-2020-II201336 (Artificial Intelligence Graduate School Support Program at UNIST) and IITP-2024-RS-2024-00360227 (Leading Generative AI Human Resources Development). We also acknowledge the 2022 Research Fund (1.220140.01) of UNIST.
@@ A.1 Human Evaluation
Response acceptability evaluation. Table 3 and Figure 8 present the evaluation guidelines and annotation interface, respectively. Human raters are given two models’ responses at once and are asked to rate each response independently by choosing one of three ratings: Acceptable (Excellent), Acceptable (Sufficient), or Not Acceptable. The order of the model responses is internally randomized at each turn to avoid potential evaluation bias.
@@ A.2 Automatic Evaluations
We use the test instructions and the LLM judge from the JustEval benchmark (Lin et al., 2024). For models without safeguards, we perform only the regular evaluation using 800 instructions. The evaluation conducted in Section 6 involves safety measures, so we also use the safety evaluation suite. The evaluation prompt can be found in their official implementation. [^6] https://github.com/Re-Align/just-eval
@@
We use the ‘weighted_alpaca_eval_gpt4_turbo’ judge from AlpacaEval 2.0 (Li et al., 2023) for the automatic preference evaluation and report length-controlled win rates (Dubois et al., 2024). The evaluation prompt can be found in their official repository. [^7] https://github.com/tatsu-lab/alpaca_eval
@@ Appendix B Experimental Setup
Response in-context learning. The simplified template of URIAL (Lin et al., 2024), URIAL-R, and zero-shot template prompt used for the evaluation can be found in Tables  45 and 6, respectively. We use urial1kv4 prompt in their official repository as a base URIAL prompt. [^9] https://github.com/Re-Align/URIAL/blob/main/urial_prompts/inst_1k_v4.txt.md Full version of URIAL-R prompt can be found in our code repository. The generation of the LLMs is truncated by the response marker of URIAL (‘‘‘).
@@ Appendix C Full Experimental Results
  • Refusal evaluation results: See Tables 11 and 12.
@@ Appendix D Model Output Examples
  • Gemma-2-9B (Riviere et al., 2024): See Tables 14 and 15 (for the model trained with refusals).
@@ Appendix E Data Examples
Appendix F RT without using IT Datasets
Theoretically, RT can also be performed using general texts that do not have paired prompts. In this experiment, we explore the RT using excerpted news articles.
Dataset. We randomly select 1,000 news passages from the CC-News dataset (Hamborg et al., 2017) and process them for use as RT training data. To establish a meaningful output distribution with news articles, we clean the excerpted sentences, enhance their readability, and adjust their tone using GPT-4o (OpenAI, 2023). The prompt and examples of the training data are provided in Tables 19 and 20.
Training and evaluation. Using this dataset, we fine-tune Llama-3.1-8B and assess its instructability with JustEval.
Results. Table 21 presents the evaluation results. The results indicate that the RT model, trained on excerpted news articles, can generate appropriate responses to various instructions, with a style reflecting the concise nature of its training data. Examples of the model’s outputs are shown in Table 22.
@@ Appendix G Refining Response Distribution
Our experiments demonstrate that establishing a response distribution alone can make LLMs instructable, thanks to the inherent abilities acquired during pre-training. In this section, we explore whether refining the training response distribution can lead to improvements in user preference for the outputs. While previous works have shown that techniques such as feedback learning or fully regenerating IT dataset responses using aligned LLMs can enhance user preferences (Bai et al., 2022; Peng et al., 2023; Ivison et al., 2024), we investigate whether refining the response distributions in IT or RT data can yield similar benefits.
@@ G.1 Experimental Setup
Response refinement. We focus on refining the fundamental characteristics of the response distribution, specifically examining three core elements that correlate with response quality: clarity, structure, and tone. To achieve this refinement, we use a strong instruction-following LLM, Llama-3.1-70B-Instruct (Dubey et al., 2024), with a manually crafted refinement prompt. The responses from the Alpaca, Dolly, and LIMA datasets are refined according to the following guidelines:
@@
The prompt and the example of refined data are available in Tables 23 and 24.
@@ G.2 Results
Refined response distribution improves user preference. As shown in Table 25, the pairwise evaluation results indicate that both IT and RT models trained with the refined responses substantially outperform their counterparts trained on the original responses. The improvements are consistent across different base models and datasets, except for Alpaca. This exception may be due to Alpaca’s responses having limited room for improvement, as they are generated using GPT-4, a highly-aligned model using human feedback. These findings suggest that the training response distribution substantially contributes to the model’s ability to generate effective outputs, emphasizing the importance of the training response distribution in yielding instructable model. The examples of model output can be found in Table 26.
@@ Appendix H Note on Potential Contamination Concerns
We use widely recognized open-weight models in our main experiments. However, since their pre-training datasets are not publicly disclosed, there may be concerns that these datasets contain text conforming to our chat template with special tokens (<|user|> and <|assistant|>). However, our main results demonstrate that base models cannot appropriately respond to instructions without an established response distribution, suggesting that such data were not meaningfully included in their pre-training datasets. Furthermore, we verify that our core observation holds for the open-data OLMo-7B model, whose authors have confirmed that such data are not included in its pre-training dataset (Hewitt et al., 2024). The results for this model are presented in Table 27.
@@
Table 4: URIAL prompt template used in our experiment (official urial1kv4 version). Exact demonstrations are omitted due to space constraints. The full prompt is available in their official repository (Lin et al., 2024).
@@ 2. Refusal:
Llama-3.1-8B (Touvron et al., 2023) - Untuned 0.05 0.00 0.94
@@
Gemma-2-9B (Riviere et al., 2024) - Untuned 0.06 0.00 0.94
@@
Mistral-7B-v0.3 (Jiang et al., 2023) - Untuned 0.04 0.00 0.96
@@
Gemma-2-2B (Riviere et al., 2024) - Untuned 0.01 0.00 0.99
@@
Table 8: Response acceptability evaluation results for IT and RT Models. The results indicate that both model types appropriately respond to a wide range of instructions.
@@
Llama-3.1-8B (Touvron et al., 2023) Alpaca IT 59.83 37.40 55.37 58.48 51.02 75.35 56.24
@@
- Untuned 63.36 33.6 60.04 66.34 55.72 80.14 59.87
Gemma-2-9B (Riviere et al., 2024) Alpaca IT 65.22 39.00 52.68 61.33 67.78 76.88 60.48
@@
- Untuned 69.04 33.8 61.09 74.42 69.9 81.28 64.92
Mistral-7B-v0.3 (Jiang et al., 2023) Alpaca IT 53.84 30.20 50.02 54.00 33.89 73.50 49.24
@@
- Untuned 59.2 33.6 60.91 64.56 40.33 80.25 56.48
Gemma-2-2B (Riviere et al., 2024) Alpaca IT 46.84 33.00 50.55 56.35 21.53 74.48 47.13
@@
- Untuned 49.34 31.2 54.95 63.53 28.73 78.4 51.03
Table 9: Core capabilities evaluation results for IT and RT Models. We observe no significant performance gap between IT and RT models.
@@
Llama-3.1-8B (Touvron et al., 2023) Alpaca IT 4.48 4.33 4.80 3.52 3.97 4.22
@@
Gemma-2-9B (Riviere et al., 2024) Alpaca IT 4.53 4.46 4.84 3.60 3.95 4.28
@@
Mistral-7B-v0.3 (Jiang et al., 2023) Alpaca IT 4.44 4.27 4.78 3.54 3.95 4.20
@@
Gemma-2-2B (Riviere et al., 2024) Alpaca IT 4.04 3.87 4.51 3.21 3.66 3.86
@@
Table 10: GPT-4 response quality evaluation results for IT and RT models. RT models perform similarly to IT models across all metrics in the JustEval benchmark  (Lin et al., 2024).
@@
Base LLM Base Dataset Method # of Mixed Safety Examples AdvBench HarmBench Malicious Instruct XSTest (unsafe) Average XSTest (benign)
@@
Gemma-2-9B (Riviere et al., 2024) Alpaca IT 0 0.29 0.13 0.20 0.66 0.32 0.07
@@
Gemma-2-2B (Riviere et al., 2024) Alpaca IT 0 0.19 0.29 0.05 0.24 0.19 0.05
@@
Table 11: Refusal evaluation results for IT and RT models (Gemma-2-9B and Gemma-2-2B) trained with the refusal examples. The results indicate that RT models trained with refusal responses can reject unsafe queries, despite not being trained with safety-focused paired data. However, we observe a noticeable gap between Gemma-2-2B IT and RT models. This gap largely diminishes as the base model size increases.
Base LLM Base Dataset Method # of Mixed Safety Examples AdvBench HarmBench Malicious Instruct XSTest (unsafe) Average XSTest (benign)
@@
Llama-3.1-8B (Dubey et al., 2024) Alpaca IT 0 0.35 0.22 0.30 0.65 0.38 0.09
@@
Mistral-7B-v0.3 (Jiang et al., 2023) Alpaca IT 0 0.17 0.20 0.08 0.36 0.20 0.06
@@
Table 12: Refusal evaluation results for IT and RT models (Llama-3.1-8B and Mistral-7B-v0.3) trained with the refusal examples. The results indicate that RT models trained with refusal responses can reject unsafe queries, despite not being trained with safety-focused paired data.
@@
Table 13: Outputs generated by Llama-3.1-8B IT and RT models trained on Alpaca. Both generated responses adequately address the given instructions.
@@
Table 14: Outputs generated by Gemma-2-9B IT and RT models trained on Alpaca. Both generated responses adequately address the given instructions.
@@
Table 15: Outputs generated by Gemma-2-9B IT and RT models trained on Alpaca mixed with prompt-refusal pairs. Both models refuse to provide the requested information.
@@
Table 16: Outputs generated by Mistral-7B-v0.3 IT and RT models trained on Dolly. Both generated responses adequately address the given instructions.
@@
Table 17: Outputs generated by Gemma-2-2B IT and RT Models trained on LIMA. Both generated responses adequately address the given instructions.

### v3

@@ 7 Conclusion
We hypothesize that the pre-training stage enables LLMs to develop the ability to process instructions. To test this, we propose Response Tuning (RT), a method that removes instructions from IT and focuses solely on learning a response distribution. Our extensive experiments demonstrate that RT models—trained only on responses without paired instructions—can effectively respond to a wide range of user queries. Moreover, we observe that they can identify and correspondingly reject unsafe requests by invoking a safety policy learned solely from response data. These observations also extend to an in-context learning setting. Such results show that establishing an adequate output distribution alone can yield instructable models, supporting our hypothesis. Taken together, our work contributes to understanding how LLMs become instructable agents and suggests the potential of extensive inherent capabilities developed during pre-training.
@@
As we discussed in Section 2 and 3, RT is designed as a verification tool—intentionally designed as a simple ablation of IT. Since IT introduces additional instruction-response mappings during fine-tuning (the effect of which is discussed in Section 4), RT exhibits limited performance compared to IT. Our work, utilizing RT rather than advocating for RT as a practical alignment method, seeks to provide insights into how LLMs shift from pre-trained models to instructable agents.
@@
Our study involves human evaluations to evaluate instruction-following LLMs. The evaluators were hired in compliance with local laws and were paid appropriate compensation. The authors manually reviewed the LLM responses flagged by the OpenAI moderation API and confirmed that these pose no harm to human evaluators. In addition, evaluators had the right to immediately stop the evaluation if they wished, and were encouraged to discuss any discomfort with the authors. While we publicly release the codes for safety evaluations, we decide not to release the refusal judge validation set to prevent potential misuse of unsafe or illegal information.
@@ Acknowledgements
We thank the reviewers for their valuable feedback. Seokhyun An is supported by the Korea Presidential Science Scholarship. This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Government of the Republic of Korea (MSIT) (RS-2020-II201336, Artificial Intelligence Graduate School Program (UNIST); RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH); and RS-2024-00360227 (Leading Generative AI Human Resources Development)) and by the 2022 Research Fund (1.220140.01) of UNIST.
@@ A.1 Human Evaluation
Response acceptability evaluation. Table 6 and Figure 8 present the evaluation guidelines and annotation interface, respectively. Human raters are given two models’ responses at once and are asked to rate each response by choosing one of three ratings: Acceptable (Excellent), Acceptable (Sufficient), or Not Acceptable. The order of the model responses is internally randomized at each turn to avoid potential evaluation bias.
@@ A.2 Automatic Evaluations
We use the test instructions and the LLM judge from the JustEval benchmark (Lin et al., 2024). For models without safeguards, we evaluate on the 800 regular instructions. The evaluations in Section 6 also use the safety subset. The evaluation prompt can be found in the official code repository. [^6] https://github.com/Re-Align/just-eval
@@
We use the ‘weighted_alpaca_eval_gpt4_turbo’ judge from AlpacaEval 2.0 (Li et al., 2023) for the automatic preference evaluation and report length-controlled win rates (Dubois et al., 2024). The evaluation prompt can be found in the official code repository. [^7] https://github.com/tatsu-lab/alpaca_eval
@@ Appendix B Experimental Setup
Response in-context learning. The simplified template of URIAL (Lin et al., 2024), URIAL-R, and zero-shot template prompt used for the evaluations can be found in Table  78 and 9, respectively. We use urial1kv4 prompt in the official code repository as a base URIAL prompt. [^9] https://github.com/Re-Align/URIAL/blob/main/urial_prompts/inst_1k_v4.txt.md Full version of URIAL-R prompt can be found in our code repository. The generation of the LLMs is truncated by the response marker of URIAL (‘‘‘).
@@ Appendix C Full Experimental Results
  • InFoBench (Qin et al., 2024) evaluation results: See Table 14.
  • Refusal evaluation results: See Table 15 and 16.
@@ Appendix D Model Output Examples
  • Gemma-2-9B (Riviere et al., 2024): See Table 18 and 19 (for the model trained with refusals).
@@ Appendix E Data Examples
Appendix F RT without IT Datasets
Theoretically, RT can be performed on general text that lacks paired prompts. Here, we explore RT with excerpts from news articles.
Dataset. We randomly sample 1,000 news passages from the CC-News dataset (Hamborg et al., 2017). To establish a meaningful output distribution with news articles, we clean the excerpts, improve readability, and adjust tone using GPT-4o (OpenAI, 2023). The prompt and illustrative training examples appear in Table 23 and 24.
Training and evaluation. We fine-tune Llama-3.1-8B on this dataset and assess instructability with JustEval.
Results. Table 25 presents the evaluation results. The RT model trained on news excerpts produces appropriate responses to diverse instructions, with a style that reflects the concision of its training data. Example outputs are shown in Table 26.
@@ Appendix G Refining Response Distribution
Our experiments demonstrate that establishing a response distribution alone can make LLMs instructable, owing to abilities acquired during pre-training. In this section, we explore whether refining the training response distribution can lead to improvements in user preference for the outputs. While previous works have shown that techniques such as feedback learning or fully regenerating responses in IT datasets with aligned LLMs can enhance user preferences (Bai et al., 2022; Peng et al., 2023; Ivison et al., 2024), we investigate whether refining the response distribution in IT or RT data yields similar benefits.
@@ G.1 Experimental Setup
Response refinement. We refine core attributes of the response distribution, focusing on three elements correlated with response quality: clarity, structure, and tone. To perform the refinement, we use a strong instruction-following LLM (Llama-3.1-70B-Instruct (Dubey et al., 2024)) with a manually crafted refinement prompt. Responses from the Alpaca, Dolly, and LIMA datasets are refined according to the following guidelines:
@@
The prompt and the example of refined data are available in Table 27 and 28.
@@ G.2 Results
Refined response distribution improves user preference. As shown in Table 29, the pairwise evaluation results indicate that both IT and RT models trained with the refined responses substantially outperform their counterparts trained on the original responses. The improvements are consistent across different base models and datasets, except for Alpaca. This exception may be due to Alpaca’s responses having limited room for improvement, as they are generated using GPT-4, a highly aligned model trained with human feedback. These findings indicate that the training response distribution contributes substantially to output quality, underscoring its importance for producing more effective models. The examples of model output can be found in Table 30.
@@ Appendix H Note on Potential Contamination Concerns
We use widely adopted open-weight models in our main experiments. However, since their pre-training datasets are not publicly disclosed, there may be concerns that they contain text that conforms to our chat template with special tokens. Our main results, however, show that base models do not respond appropriately to instructions without an established response distribution, suggesting that such data are not materially present in their pre-training data. In addition, we validate our core observation on OLMo-7B (Groeneveld et al., 2024), an open-data model whose authors report that no such data were included in its pre-training dataset (Hewitt et al., 2024). Results for this model appear in Table 31.
@@
Table 7: URIAL prompt template used in our experiment (official urial1kv4 version). Exact demonstrations are omitted due to space constraints. The full prompt is available in Lin et al. (2024).
@@ 2. Refusal:
Llama-3.1-8B
(Touvron et al., 2023)
- Untuned 0.05 0.00 0.94
@@
Gemma-2-9B
(Riviere et al., 2024)
- Untuned 0.06 0.00 0.94
@@
Mistral-7B-v0.3
(Jiang et al., 2023)
- Untuned 0.04 0.00 0.96
@@
Gemma-2-2B
(Riviere et al., 2024)
- Untuned 0.01 0.00 0.99
@@
Table 11: Response acceptability evaluation results for RT and IT models. The results indicate that both model types appropriately respond to a wide range of instructions.
@@
Llama-3.1-8B
(Touvron et al., 2023)
Alpaca IT 59.83 37.40 55.37 58.48 51.02 75.35 56.24
@@
- Untuned 63.36 33.60 60.04 66.34 55.72 80.14 59.87
Gemma-2-9B
(Riviere et al., 2024)
Alpaca IT 65.22 39.00 52.68 61.33 67.78 76.88 60.48
@@
- Untuned 69.04 33.80 61.09 74.42 69.90 81.28 64.92
Mistral-7B-v0.3
(Jiang et al., 2023)
Alpaca IT 53.84 30.20 50.02 54.00 33.89 73.50 49.24
@@
- Untuned 59.20 33.60 60.91 64.56 40.33 80.25 56.48
Gemma-2-2B
(Riviere et al., 2024)
Alpaca IT 46.84 33.00 50.55 56.35 21.53 74.48 47.13
@@
- Untuned 49.34 31.20 54.95 63.53 28.73 78.40 51.03
Table 12: Core capabilities evaluation results for RT and IT models. We observe no significant performance gap between IT and RT models.
@@
Llama-3.1-8B
(Touvron et al., 2023)
Alpaca IT 4.48 4.33 4.80 3.52 3.97 4.22
@@
Gemma-2-9B
(Riviere et al., 2024)
Alpaca IT 4.53 4.46 4.84 3.60 3.95 4.28
@@
Mistral-7B-v0.3
(Jiang et al., 2023)
Alpaca IT 4.44 4.27 4.78 3.54 3.95 4.20
@@
Gemma-2-2B
(Riviere et al., 2024)
Alpaca IT 4.04 3.87 4.51 3.21 3.66 3.86
@@
Table 13: GPT-4 response quality evaluation results for RT and IT models. RT models perform similarly to IT models across all metrics in the JustEval benchmark  (Lin et al., 2024).
@@
Model InFoBench
Llama-3.1-8B
++Alpaca
IT 0.77
RT 0.69
Untuned 0.36
Gemma-2-9B
++Alpaca
IT 0.79
RT 0.74
Untuned 0.28
Mistral-7B-v0.3
++Alpaca
IT 0.75
RT 0.71
Untuned 0.34
Gemma-2-2B
++Alpaca
IT 0.68
RT 0.60
Untuned 0.21
Table 14: Average DRFR for RT and IT models. The models are evaluated on 500 decomposable test instructions from InFoBench (Qin et al., 2024). The results indicate that the models trained with paired examples more sensitively capture prompt requirements.
Base LLM
Base
Dataset
Method
# of Mixed
Safety
Examples
AdvBench HarmBench Malicious Instruct XSTest (unsafe) Average XSTest (benign)
@@
Gemma-2-9B
(Riviere et al., 2024)
Alpaca IT 0 0.29 0.13 0.20 0.66 0.32 0.07
@@
Gemma-2-2B
(Riviere et al., 2024)
Alpaca IT 0 0.19 0.29 0.05 0.24 0.19 0.05
@@
Table 15: Refusal evaluation results for RT and IT models (Gemma-2-9B and Gemma-2-2B) trained with the refusal examples. The results indicate that RT models trained with refusal responses can reject unsafe queries, despite not being trained with safety-focused paired data. However, we observe a noticeable gap between Gemma-2-2B IT and RT models. This gap largely diminishes as the base model size increases.
Base LLM
Base
Dataset
Method
# of Mixed
Safety
Examples
AdvBench HarmBench Malicious Instruct XSTest (unsafe) Average XSTest (benign)
@@
Llama-3.1-8B
(Dubey et al., 2024)
Alpaca IT 0 0.35 0.22 0.30 0.65 0.38 0.09
@@
Mistral-7B-v0.3
(Jiang et al., 2023)
Alpaca IT 0 0.17 0.20 0.08 0.36 0.20 0.06
@@
Table 16: Refusal evaluation results for RT and IT models (Llama-3.1-8B and Mistral-7B-v0.3) trained with the refusal examples. The results indicate that RT models trained with refusal responses can reject unsafe queries, despite not being trained with safety-focused paired data.
@@
Table 17: Outputs generated by Llama-3.1-8B RT and IT models trained on Alpaca. Both generated responses adequately address the given instructions.
@@
Table 18: Outputs generated by Gemma-2-9B RT and IT models trained on Alpaca. Both generated responses adequately address the given instructions.
@@
Table 19: Outputs generated by Gemma-2-9B RT and IT models trained on Alpaca mixed with prompt-refusal pairs. Both models refuse to provide the requested information.
@@
Table 20: Outputs generated by Mistral-7B-v0.3 RT and IT models trained on Dolly. Both generated responses adequately address the given instructions.
@@
Table 21: Outputs generated by Gemma-2-2B RT and IT models trained on LIMA. Both generated responses adequately address the given instructions.